OpenStudy (anonymous):

how do you solve 4*x^2 + 49 = 0

OpenStudy (anonymous):

\[4x^{2} + 49 =0\]

OpenStudy (anonymous):

I'm sure there are other ways, but you can do: 4x^2 + 49 = 0 4x^2 = -49 x^2 = -49 / 4 x = sqrt ( -49 / 4) which results in a complex solution. There is probably a nicer way to do it.

OpenStudy (anonymous):

4x^2+49=0 4x^2=-49 4x^2=49i^2 x^2=(49/4)*i^2 x=(7/2)i complex number for i

OpenStudy (anonymous):

after i subtract 49 to both sides do i devide by 4 first? or square root?

OpenStudy (anonymous):

Divide by 4 first

OpenStudy (anonymous):

im left with x\[x ^{2} = 49\div4\]

OpenStudy (anonymous):

you left the minus sign

OpenStudy (anonymous):

-49

OpenStudy (anonymous):

correct (except for the minus sign), now take the square root of both sides, which solves for x so you get x =sqrt ( -49/4 )

OpenStudy (anonymous):

should be x^2= -49/4 instead

OpenStudy (anonymous):

after that how would i square root?

OpenStudy (anonymous):

Notice that 49 is 7^2 and 4 is 2^2, so sqrt ( -49 / 4) = 7/2, however you are taking the square root of a negative number, so you need to multiply by a complex factor i. So sqrt ( -49 / 4) = 7/2i, to be precise

OpenStudy (anonymous):

i need imaginary numbers if any

OpenStudy (anonymous):

i^2 = -1

OpenStudy (anonymous):

the answer would be \[\pm7i/2?\]

OpenStudy (anonymous):

it does not need to be plus or minus, otherwise the answer is correct

OpenStudy (anonymous):

the i takes care of the plus or minus

OpenStudy (anonymous):

but when u squareroot an equation isn't there always plus or minus?

OpenStudy (anonymous):

the answer is just imaginary.... not a real number.

OpenStudy (anonymous):

If you are going to include i, you do not need to have a plus or minus attached to the square root. To do so would be redundant. If you take the square root of -1, you get i. So the square root of -49, is 7i. It actually doesn't make sense to say sqrt (-1) = +- i

OpenStudy (anonymous):

:)